OpenStudy (anonymous):

Compute a value for y that satisfies the equation below. –3y – 12 = –4

OpenStudy (dmndlife24):

To do that we have to first solve for y... -3y -12 = -4 Add 12 to both sides... -3y = 8 divide by -3 on both sides to isolate y... y = -8/3 Check to see if it satisfies the equation by plugging (-8/3) for y in the original equation... -3(-8/3) - 12 = -4 8 - 12 = -4 -4 = -4 True It checks so the solution set is {-8/3}